# Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B

### v1

@@ 4.2 Brightness Comparison to GRB sample and physical scenarios for GRB 230204B
The optical luminosity of GRB 230204B (R/r-band and g-band data) has been compared with the population of GRBs with measured redshifts (Dainotti et al., 2024), providing valuable insights into its optical emission characteristics. The temporal evolution of GRB 230204B’s optical luminosity (see Figure 9) reveals an exceptionally bright initial emission, followed by a decay phase consistent with a forward shock model in a homogeneous interstellar medium (ISM). This places it among the brightest optical GRBs ever observed at the epoch of its detection, making it a standout event in the context of optical afterglow studies. The extreme optical brightness is an order of magnitude higher than other highly energetic bursts ( 1.3 ks post-burst) such as GRB 080319B (Racusin et al., 2008; Pandey et al., 2009), GRB 990123 (Castro-Tirado et al., 1999), GRB 190114C (Misra et al., 2021; Gupta et al., 2021a), and GRB 221009A. Furthermore, we have also examined the optical luminosity of GRB 230204B at the epoch of its detection as a function of redshift for a population of GRBs with measured redshifts (Dainotti et al., 2024). We noted that GRB 230204B has the highest optical luminosity ( 1.1-2.4 ks post-burst) with respect to all known GRBs with measured redshifts at the same epoch (see Figure 9). This suggests that GRB 230204B could represent a highly efficient energy transfer to the optical afterglow, or it might be related to the properties of its surrounding environment, jet structure, or progenitor system, contributing to its extraordinary luminosity.
@@
Additionally, we noted that GRB 230204B follows an anti-correlation between the optical decay index and optical luminosity (Oates et al., 2012, 2015). There are several scenarios where such a correlation could be explained. Continued energy injection from a central engine, such as a magnetar or a rapidly spinning black hole, could initially produce a very bright afterglow (Zhang & Mészáros, 2001). However, if the injection is short-lived or if the engine rapidly loses power, the afterglow could exhibit a steep decay as the additional energy source wanes (Zhang & Mészáros, 2001; Dai & Lu, 1998). Our afterglow analysis did not show evidence of central engine activity (flare, plateau, variability, etc). A reverse shock could also produce a very luminous optical flash (Zhang & Kobayashi, 2005). If the reverse shock is strong, it can dominate the early afterglow. Once the reverse shock fades, the light curve might steepen significantly, leading to a rapid decay. For GRB 230204B, however, our afterglow modeling suggests that a forward shock model could explain the early optical emission. Another possibility involves a structured jet, where a bright core with steep wings might cause a steep decay if the emission transitions from being core-dominated to wing-dominated, or if there is a slight change in the viewing angle (Rossi et al., 2002; Granot & Kumar, 2003). However, our afterglow modeling suggests that a top-hat jet could explain the afterglow of GRB 230204B, so we discarded the structured jet for this burst. If the GRB is in a wind-like medium (density r-2), the early afterglow can be very bright (Chevalier & Li, 2000). A change in the density profile (e.g., transition to a uniform medium) could cause the afterglow to decay steeply (Panaitescu & Kumar, 2002; Starling et al., 2008). However, for GRB 230204B, our afterglow modeling suggests that an ISM-like medium could explain the broadband emission. A jet break occurs when the relativistic beaming angle of the jet becomes wider than the observer’s line of sight (Racusin et al., 2009). This typically results in a steepening of the afterglow light curve (Liang et al., 2008). An achromatic temporal slope of -2 at late times could indicate that the break occurred, however, for GRB 230204B, synchronous broadband data is not available to confirm such behavior, also such an early jet break is not expected so we discarded the possibility. GRB afterglows depend strongly on the observer’s viewing angle relative to the jet’s core (Granot et al., 1999; van Eerten & MacFadyen, 2013). When an observer views the GRB slightly off-axis but still close to the jet core, the observed afterglow appears very bright initially. This is due to the deceleration of the relativistic jet as it interacts with the surrounding medium, causing the beamed emission to broaden and spread into the observer’s line of sight. As the jet slows further and its emission becomes more isotropic, the light curve steepens, showing a rapid decline (Ryan et al., 2015). Such behavior suggests that GRB 230204B is observed close to the jet axis, as indicated by its narrow inferred viewing angle from afterglow modeling. The steep early optical decay and extreme luminosity support this interpretation.
@@
In Figure 12, we present a comparative analysis of the X-ray luminosity light curves of GRB 230204B with a sample of other GRBs that have measured redshifts (Gupta et al., 2024b). The X-ray flux data for these GRBs were sourced from the Swift XRT light curve repository444https://www.swift.ac.uk. To enable a direct comparison, the light curves are presented in the rest frame in the k-corrected co-moving bandpass, allowing for a consistent examination of intrinsic luminosity evolution across the sample. GRB 230204B is characterized by limited data, with only two available data points recorded at a late epoch, depicted in red on the plot. Despite the scarcity of data, the observed luminosity for GRB 230204B during this late phase is notably consistent with the average luminosity values observed in the comparative GRB sample at a similar epoch. This suggests that GRB 230204B’s X-ray luminosity profile falls within the expected range for GRBs in this phase.
@@ 5 SUMMARY
Furthermore, the afterglow data collected across radio, optical, and X-ray wavelengths showed that GRB 230204B’s luminosity at various stages aligns with typical long-duration GRBs. Optical follow-up using multiple ground-based observatories revealed a rapid, early afterglow decay rate, which was more luminous than most GRBs at a similar epoch. Its high optical luminosity and bright afterglow suggest efficient energy transfer from the burst to the surrounding medium, or an environment conducive to high luminosity, possibly due to a structured jet.
@@
Overall, our analysis of GRB 230204B underscores its importance as a reference for studying the emission mechanisms, jet structures, and afterglow behaviors of long-duration GRBs. The findings contribute to the growing understanding of GRB diversity and emphasize the role of broadband follow-up observations in exploring the complex evolution of these powerful cosmic events. Future studies of similar events, using well-coordinated observations across the spectrum, will be crucial in further refining our understanding of GRB physics and jet composition.

### v2

@@ 4.2 Brightness Comparison to GRB sample and physical scenarios for GRB 230204B
In this Section, we compared the optical, X-ray, and radio afterglow brightness of GRB 230204B with a larger population of GRBs with measured redshifts and explored the physical scenarios.
@@
The optical luminosity of GRB 230204B (R/r-band and g-band data) has been compared with the population of GRBs with measured redshifts (2024MNRAS.533.4023D), providing valuable insights into its optical emission characteristics. The temporal evolution of GRB 230204B’s optical luminosity (see Figure 10) reveals an exceptionally bright initial emission, followed by a decay phase consistent with a forward shock model in a homogeneous interstellar medium (ISM). This places it among the brightest optical GRBs ever observed at the epoch of its detection, making it a standout event in the context of optical afterglow studies. The extreme optical brightness is an order of magnitude higher than other highly energetic bursts ( 1.3 ks post-burst) such as GRB 080319B (2008Natur.455..183R; 2009A&A...504...45P), GRB 990123 (1999Sci...283.2069C), GRB 190114C (2021MNRAS.504.5685M; 2021RMxAC..53..113G), and GRB 221009A. Furthermore, we have also examined the optical luminosity of GRB 230204B at the epoch of its detection as a function of redshift for a population of GRBs with measured redshifts (2024MNRAS.533.4023D). We noted that GRB 230204B has the highest optical luminosity ( 1.1-2.4 ks post-burst) with respect to all known GRBs with measured redshifts at the same epoch (see Figure 10). This suggests that GRB 230204B could represent a highly efficient energy transfer to the optical afterglow, or it might be related to the properties of its surrounding environment, jet structure, or progenitor system, contributing to its extraordinary luminosity.
@@
Additionally, we noted that GRB 230204B follows an anti-correlation between the optical decay index and optical luminosity (2012MNRAS.426L..86O; 2015MNRAS.453.4121O). There are several scenarios where such a correlation could be explained. Continued energy injection from a central engine, such as a magnetar or a rapidly spinning black hole, could initially produce a very bright afterglow (2001ApJ...552L..35Z). However, if the injection is short-lived or if the engine rapidly loses power, the afterglow could exhibit a steep decay as the additional energy source wanes (2001ApJ...552L..35Z; 1998A&A...333L..87D). Our afterglow analysis did not show evidence of central engine activity (flare, plateau, variability, etc). A reverse shock could also produce a very luminous optical flash (2005ApJ...628..315Z). If the reverse shock is strong, it can dominate the early afterglow. Once the reverse shock fades, the light curve might steepen significantly, leading to a rapid decay. For GRB 230204B, however, our afterglow modeling suggests that a forward shock model could explain the early optical emission. Another possibility involves a structured jet, where a bright core with steep wings might cause a steep decay if the emission transitions from being core-dominated to wing-dominated, or if there is a slight change in the viewing angle (2002MNRAS.332..945R; 2003ApJ...591.1086G). However, our afterglow modeling suggests that a top-hat jet could explain the afterglow of GRB 230204B, so we discarded the structured jet for this burst. If the GRB is in a wind-like medium (density r-2), the early afterglow can be very bright (2000ApJ...536..195C). A change in the density profile (e.g., a transition from a wind to a uniform medium) can also produce a steep decay in the afterglow light curve (2002ApJ...571..779P; 2008ApJ...672..433S). For GRB 230204B, we modeled the afterglow in both ISM and wind-like environments. While the wind scenario can reproduce some features of the light curve, our broadband analysis shows that the ISM-like medium provides a better overall fit to the data. A jet break occurs when the relativistic beaming angle of the jet becomes wider than the observer’s line of sight (2009ApJ...698...43R). This typically results in a steepening of the afterglow light curve (2008ApJ...675..528L). An achromatic temporal slope of -2 at late times could indicate that the break occurred; however, for GRB 230204B, synchronous broadband data is not available to confirm such behavior, and such an early jet break is not expected, so we discarded the possibility. GRB afterglows depend strongly on the observer’s viewing angle relative to the jet’s core (1999ApJ...513..679G; 2013ApJ...767..141V). When an observer views the GRB slightly off-axis but still close to the jet core, the observed afterglow appears very bright initially. This is due to the deceleration of the relativistic jet as it interacts with the surrounding medium, causing the beamed emission to broaden and spread into the observer’s line of sight. As the jet slows further and its emission becomes more isotropic, the light curve steepens, producing a rapid decline (2015ApJ...799....3R). For GRB 230204B, our afterglow modeling yields an observer angle of rad and a jet core angle of rad. Although the nominal values of the observer and jet core angles suggest that the line of sight is not exactly aligned with the jet axis. However, within the reported uncertainties, the two angles are consistent. This is possibly compatible with the steep early optical decay and high luminosity.
@@
In Figure 15, we present a comparative analysis of the X-ray luminosity light curves of GRB 230204B with a sample of other GRBs that have measured redshifts (2024arXiv240904871G). The X-ray flux data for these GRBs were sourced from the Swift XRT light curve repository444https://www.swift.ac.uk. To enable a direct comparison, the light curves are presented in the rest frame in the k-corrected co-moving bandpass, allowing for a consistent examination of intrinsic luminosity evolution across the sample. GRB 230204B is characterized by limited data, with only two available data points recorded at a late epoch, depicted in red on the plot. Despite the scarcity of data, the observed luminosity for GRB 230204B during this late phase is notably consistent with the average luminosity values observed in the comparative GRB sample at a similar epoch. This suggests that GRB 230204B’s X-ray luminosity profile falls within the expected range for GRBs in this phase.
@@ 5 SUMMARY
Furthermore, the afterglow data collected across radio, optical, and X-ray wavelengths showed that GRB 230204B’s luminosity at various stages aligns with typical long-duration GRBs. Optical follow-up using multiple ground-based observatories (mainly using the MASTER and BOOTES robotic telescopes) revealed a rapid, early afterglow decay rate, which was more luminous than most GRBs at a similar epoch. The contributions of the MASTER and BOOTES robotic telescope networks were indispensable for the early detection and characterization of the optical afterglow of GRB 230204B. The scientific impact of the early robotic observations extends beyond simple discovery and can be understood through these key aspects:
The continuous monitoring by the MASTER and BOOTES networks over the first few hours post-burst revealed a smooth power-law decay without evidence of optical flare, reverse shock emission, or other early-time features that have been observed in other bright GRBs (e.g., GRB 080319B 2008Natur.455..183R, GRB 990123, 1999Natur.398..400A). This forward-shock-dominated evolution, established through early observations, allowing robust constraints on the microphysical parameters, including the circumburst density, jet viewing angle, jet opening angle, and magnetic field properties through broadband modeling (see Section 3.4). These parameters are fundamental to understanding the energetics and collimation of this exceptionally luminous GRB, underscoring the importance of robotic networks in the study of transient astronomical phenomena. In particular, without these rapid robotic measurements, the steep optical decay and its placement on luminosity–decay correlations would remain unconstrained. The high optical luminosity and bright afterglow of GRB 230204B indicate efficient energy transfer from the burst to the surrounding medium or an environment conducive to high luminosity. Our afterglow modeling suggests that this can be explained within the framework of a top-hat jet.
@@
Overall, our analysis of GRB 230204B highlights its significance as a benchmark for studying the emission mechanisms, and afterglow evolution of long-duration GRBs. The results not only enrich the broader understanding of GRB diversity but also underscore the indispensable role of broadband, multi-wavelength follow-up observations in probing the complex physical processes governing these events. In particular, this case study demonstrates that the value of robotic telescope networks extends far beyond rapid discovery: their early-time, high-cadence data provide essential constraints on jet dynamics, circumburst environments, and emission mechanisms that would otherwise remain ambiguous. As next-generation and expanded robotic networks become operational, their systematic ability to capture the earliest phases of GRB afterglows will be crucial for refining jet models, improving physical parameter estimates, and advancing our overall understanding of GRB physics and relativistic jet composition.
@@ Appendix A Figures
| 27405.6 | Clear | 19.6 | – | BOOTES | This work |
| 36280.3 | 17.7 | – | BOOTES | This work | |
| 37487.6 | Clear | 19.7 | – | BOOTES | This work |
| 37558.0 | 19.4 | – | BOOTES | This work | |
| 72995.8 | 19.3 | – | BOOTES | This work | |
| 74544.8 | 19.4 | – | BOOTES | This work | |